louisa is 3/4 as tall as Carla. The difference between their heights is .4 meters. How tall is each girl?
Thank you

Answer by tutorcecilia(2152)

Louisa is 3/4 as tall as Carla
Translation:
Louisa = (3/4)(Carla)=(3/4)(C)
Carla = (C)
.
The difference between their heights is .4 meters
Translation:
Louisa-Carla=.4
or
(3/4)(C)-(C)=0.4 [Simplify]
(1/4)C=0.4 [Solve for C]
C=1.6

Check by plugging all of the values back into the original equation:
(3/4)C-C=0.4
(3/4)(1.6)-(1.6)=0.4